Assume that the potential of a -0.390 mC point charge at infinity is exactly zero volts. A proton is moved along a line from 5.8 m away from the point charge to 2.3 m away. Assume the two charges are isolated.

A) Through what potential difference did the proton move?

  1. What was the change in electrical PE for the proton?
Answer #1

a)

V = k q ( 1/r1 - 1/r2)

V = 10^9* (- 0.39* 10^-3)* ( 1 / 2.3 - 1/5.8)

V = - 1.023* 10^5 V

======

b)

U = qV = 1.6* 10^-19* 1.023* 10^5 = - 1.637* 10^-14 J

=====

Comment before rate in case any doubt, will reply for sure.. goodluck